OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (mayaal):

Use y=mx+b

OpenStudy (mayaal):

@dacar are u there?

OpenStudy (anonymous):

Yeah I'm here

OpenStudy (anonymous):

I got 4x + 4y = 0

OpenStudy (anonymous):

Are you there?????

OpenStudy (mayaal):

yup

OpenStudy (mayaal):

now try to change it in the form of y=mx+b

OpenStudy (mayaal):

@ dacar?????????

OpenStudy (anonymous):

Yes I'm here

OpenStudy (mayaal):

did u try to change it

OpenStudy (anonymous):

The problem is the answer is supposed to be X _ _ =

OpenStudy (mayaal):

what do you mean?

OpenStudy (mayaal):

X _ _ =?

OpenStudy (mayaal):

like do u want to solve it for x?

OpenStudy (anonymous):

I mean the answer needs to start with X

OpenStudy (mayaal):

if it starts with x then it would be 1

OpenStudy (mayaal):

because 4/4 is 1

OpenStudy (anonymous):

Could it possibly be X + Y = -4?

OpenStudy (mayaal):

yah it can be

OpenStudy (anonymous):

Then would it be X + Y = 0?

OpenStudy (anonymous):

...

OpenStudy (mayaal):

yup i am pretty sure

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (mayaal):

ur welcome i hope its correct...